Shift report 20011006 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Oct 06 2001 - 08:20:30 EDT

  • Next message: Apache: "Shift report 20011006 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Oct  6 08:20:30 2001, Shift: 00:00-08:00
    Supervisor: JH Lee, with: 
    Log book#9 pages 103 - 105
    
    Shift Summary:
    ------------------------------------------------------------
    Beam:
      - Cogged until 2:00 and A new store at 7:30
      - AGS power supply fault between stores (replacement)
    
    Data:
      - MRS at 35degree 4A
      - FS at  12/3degree 1/3B (d3,d4=0)
      - Physics Runs: 5186,5187,5193
    
    Some Pictures: 
      "Tracks" are nicely shown in T4 (from run5187)
      *Plots for "total number of hits" in T4-1,2,3 at 3degree (d3,D4:off) 
      - http://pii3.brahms.bnl.gov/~operator/pictures/t4_1_num_hit_run5187.ps
      - http://pii3.brahms.bnl.gov/~operator/pictures/t4_2_num_hit_run5187.ps
      - http://pii3.brahms.bnl.gov/~operator/pictures/t4_3_num_hit_run5187.ps
    
    Note:
      - T3 Anode are set at 40V lower than default
    
    No new detector problems seen:
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Oct 06 2001 - 08:21:10 EDT